graph: y=4x+80, y=1/2x+10
where do they intercept?

Found 2 solutions by jorel1380, rfer:
Answer by jorel1380(2518)
Answer by rfer(12657)

1/2x+10=4x+80
-3.5x=70
x=-20
-------------
y=4*-20+80
y=-80+80
y=0
-----------
intercept (-20,0)
----------------
0=1/5*-20+10